What is the freezing point of a solution that contains 610g of benzoic acid, HC7H5O2, in125g of camphor? kf camphor =
40
1
Expert's answer
2021-09-22T00:26:54-0400


Δt = i Kf m


Molar mass of benzoic = 122.12

610/122.12= 5moles

Molar mass of camphor = 152.23

= 125/152.23 = 0.82

= 40 /5.082

= 6.87°C